Woodsman Company sells a product for $155 per unit. The variable cost is $70 per unit, and fixed costs are $408,000.

Determine (a) the break-even point in sales units and (b) the sales units required for the company to achieve a target profit of $114,240.

a. Break-even point in sales units    units
b. Break-even point in sales units required for the company to achieve a target profit of $114,240 units
Answer #1

Contribution margin=Sales-Variable cost  

=(155-70)=$85 per unit

a.Breakeven=Fixed expenses/Contribution margin

=408,000/85

=4800 units

b.Target contribution margin=Fixed cost+Target profits

=(408,000+114240)=$522240

Hence unit sales required=Target contribution margin/contribution margin per unit

=522240/85

=6144 units